OpenStudy (anonymous):

solve 8 to the power of 5 over 8 to the power of 3 show work

OpenStudy (marigirl):

think of it like \[\frac{ 8\times8\times8\times8\times8 }{ 8\times8\times8 }\]

OpenStudy (marigirl):

you can cancel out 3 of the 8 at the top from 3 of the 8 at the bottom right? and now your left with two 8 at the top..... can you see this with the powwers? what can we do the powers to end up with 8^2

OpenStudy (anonymous):

that's what i tried but im still confused

OpenStudy (anonymous):

16

OpenStudy (anonymous):

you end up with 16

OpenStudy (marigirl):

lets leave it in power form \[\frac{ 8^5 }{ 8^3 } = 8^2\]

OpenStudy (marigirl):

i think u mean 64 ..

OpenStudy (jdoe0001):

\(\bf \cfrac{\cancel{ 8}\times\cancel{ 8}\times\cancel{ 8}\times8\times8 }{ \cancel{ 8}\times \cancel{ 8}\times \cancel{ 8}}\implies ?\)

OpenStudy (anonymous):

ya 64

OpenStudy (anonymous):

how would i show my work